Judy's Boutique just paid an annual dividend of $2.65 on its common stock. The firm increases its dividend by 3.40 percent annually. What is the rate of return on this stock if the current stock price is $39.64 a share?
Answer #1
Price = Dividend/ ke-g
D1= dividend at year 1 = 2.65*103.4%
=$2.7401
Ke = Required rate of return
g = Growth Rate
39.64= 2.7401/Ke-.034
Ke -0.034= 2.7401/39.64
Ke = 0.06912+0.034
Ke 10.31%
The required rate of return is 10.31%